OpenStudy (anonymous):

Pies are ordered from the Pizza Palace according to their diameter. Large:16 inches Medium: 14 inches Small:12 inches How many more square inches are in the large pizza than in the small pizza?

OpenStudy (anonymous):

@bahrom7893 help in this one plz

OpenStudy (bahrom7893):

Large pizza = pi*r^2 = pi*(d/2)^2 = pi*(8)^2 = 64pi

OpenStudy (bahrom7893):

Small: 36pi

OpenStudy (bahrom7893):

64pi-36pi = 28pi in^2 is the difference

OpenStudy (anonymous):

so the answer is 28

OpenStudy (anonymous):

thats not in the multiple choice @bahrom7893

OpenStudy (anonymous):

plz need help

OpenStudy (bahrom7893):

28*pi = 28*3.14 = 87.92 = 88

OpenStudy (anonymous):

if the answer is not in the multiple choice u can put letter E = error